OpenStudy (anonymous):

Solve the following system of equations. 1/2x + 1/3y = 7 1/4x + 2/3y = 6

OpenStudy (anonymous):

x²+y² = 9 Subtract y² from both sides: x² = 9-y²...(1) x²-3y = 9...(2) Substitute the value of x² from (1) into (2): 9-y²-3y = 9 Subtract 9 from both sides: y²+3y = 0 y(y+3) = 0 Either: y = 0 or, y+3 = 0 >> y = -3. When y = 0, x²+0 = 9 x² = 9 x = ­­+/-3 When y = -3, x²+(-3)² = 9 x²+9 = 9 x² = 0 x = 0 Therefore the solutions are: (x,y) = (+/-3, 0) or (0,-3) got it from yahoo answers

OpenStudy (anonymous):

no one of these answer (1/6, 14) (6, 12) (10 2/3, 5) (2 4/9, -1/6)

OpenStudy (anonymous):

ok I got B

OpenStudy (anonymous):

so you did the work and you got B]